OpenStudy (anonymous):

through (2,-3) and parallel to x-5y=10

OpenStudy (anonymous):

Luigi :3

OpenStudy (luigi0210):

Get the equation in slope intercept form, find the slope, m, then use \(\Large y-y_{1}=m(x-x_{1})\) \[\Large \color{green}{Batman~}\]

OpenStudy (anonymous):

so i would set it up as y+3=1/5(x-2)?

OpenStudy (anonymous):



OpenStudy (anonymous):

if you replace x =2, y =-3 into the left hand side of the expression, what do you have?

OpenStudy (jdoe0001):

*cough* what's the slope of " x-5y=10 " ?

OpenStudy (anonymous):

-1/5x is the slope

OpenStudy (anonymous):

I know my method is weird, but it is 100% correct for this kind of problem 2-5*(-3) = 17 then x-5y = 17 is the required line.

OpenStudy (jdoe0001):

so... what would it be if we solve "x-5y=10" for y?

OpenStudy (anonymous):

y=1/5x-2

OpenStudy (jdoe0001):

so our slope will be \(\bf y={\color{brown}{ \cfrac{1}{5}}}x-2\)

OpenStudy (anonymous):

yes

OpenStudy (jdoe0001):

so a parallel line will have the same slope as that one, or 1/5 so you're really being asked to find the equation of a line with slope of 1/5 and with point of (2,-3) \(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 2}}\quad ,&{\color{blue}{ -3}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{1}{5} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

y+3=1/5(x-2)

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

so when the distributive property is applied then it would be: \[y+3=1/5x-2/5\]

OpenStudy (jdoe0001):

yeap